# Observer-robust energy condition verification for warp drive spacetimes

### 03 Mar 2026 v2

@@ 2.4 Observer optimization
where $m_{\text{flux}}=-g_{ab}\,j^{a}j^{b}$ (positive when the energy flux $j^{a}=-T^{a}{}_{b}\,u^{b}$ is causal), $m_{\text{future}}=-j_{a}n^{a}$ (positive when $j^{a}$ is future-directed, with $n^{a}$ the future-pointing unit normal), and $m_{\text{WEC}}=T_{ab}\,u^{a}u^{b}$ (positive when the energy density is non-negative). A negative value of any component indicates a DEC violation for that observer. The three-term minimum in equation (17) is piecewise smooth but not $C^{1}$ at branch-switching boundaries; gradients exist almost everywhere, and the multi-start BFGS optimizer handles this reliably in practice. A log-sum-exp soft-minimum approximation could be employed if smoother gradients are needed. An alternative design would run three separate optimizations (one per DEC sub-condition) and take the minimum of the minima; this would avoid the unit-mixing landscape and improve interpretability of the “worst-case observer” at some additional computational cost. Because $m_{\text{WEC}}$ has dimensions of energy density while $m_{\text{flux}}$ has dimensions of (energy density)2, we do not combine them into a single margin for severity purposes. For violation detection, the DEC is violated whenever $m_{\text{DEC}}^{*}<0$ (equation 17); the sign logic is correct regardless of units. For Type I stress-energy with $\rho\geq 0$ and $\rho\geq|p_{i}|$, the algebraic condition (equation 13) guarantees that $j^{a}$ is causal and future-directed for every future-directed timelike $u^{a}$, making the three-term check redundant; the algebraic slack $\min_{i}(\rho-|p_{i}|)$ then serves as the primary DEC result with homogeneous dimensions (energy density). At non-Type I points, the optimizer evaluates all three terms of equation (17) simultaneously (Section 8). For the majority of grid points, which classify as Type I, the algebraic slack is the primary DEC result; at non-Type I points the capped extremum is used instead.
@@
The optimization uses BFGS [13] implemented in Optimistix [14], a pure-JAX optimization library. For timelike observers (WEC/SEC/DEC), the optimizer works over an unconstrained 3-vector $\mathbf{w}\in\mathbb{R}^{3}$, with the rapidity $\zeta=|\mathbf{w}|$ and boost direction $\hat{s}=\mathbf{w}/|\mathbf{w}|$. This parameterization naturally includes the Eulerian observer at $\mathbf{w}=0$ (the removable singularity $\mathbf{w}/|\mathbf{w}|$ at the origin is handled via the smooth approximation $|\mathbf{w}|_{\varepsilon}=\sqrt{\mathbf{w}\cdot\mathbf{w}+\varepsilon^{2}}$ with $\varepsilon=10^{-12}$, which is $C^{\infty}$ everywhere including $\mathbf{w}=0$ and avoids discontinuous gradients that could confuse quasi-Newton Hessian updates near the Eulerian point; see Section 8). This parameterization avoids angular coordinate singularities and ensures a smooth optimization landscape. A soft cap $|\mathbf{w}|\mapsto\zeta_{\max}\tanh(|\mathbf{w}|/\zeta_{\max})$ bounds the rapidity at $\zeta_{\max}=5$ (corresponding to $\gamma=\cosh 5\approx 74$); the $\tanh$ map is chosen because it smoothly saturates with unit gradient near the origin, preserving optimizer sensitivity for moderate boosts while enforcing a strict upper bound. For the NEC, the null direction is parameterized via stereographic projection: a 2-vector $\mathbf{w}\in\mathbb{R}^{2}$ maps to a direction on $S^{2}$ via $\hat{s}=(2w_{1},\,2w_{2},\,1-|\mathbf{w}|^{2})/(1+|\mathbf{w}|^{2})$, then $k^{a}=e_{0}^{a}+\hat{s}^{i}\,e_{i}^{a}$. This covers all of $S^{2}$ smoothly without polar coordinate singularities ($\mathbf{w}=0\to$ north pole; $|\mathbf{w}|\to\infty\to$ south pole). Because the south pole requires $|\mathbf{w}|\to\infty$, optimizer convergence near it can be slow; the multi-start pool includes near-pole initializations (axis-aligned directions with large $|\mathbf{w}|$) to mitigate this.
@@ 2.5 Geodesic integration
which is integrated simultaneously with the geodesic equation as a coupled 16-component ODE system $y=(x^{\mu},v^{\mu},\xi^{\mu},w^{\mu})$, where $v^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}\lambda$ and $w^{\mu}=D\xi^{\mu}/\mathrm{d}\lambda$. The explicit first-order form is
@@ 3 The warpax toolkit
warpax is a Python package built on JAX [6] for hardware-accelerated numerical computation with automatic differentiation. The toolkit is organized into six computational modules:
@@
Each metric is an Equinox module implementing a pointwise signature $x^{\mu}\mapsto g_{ab}(x)$, where $x^{\mu}$ is a 4-vector of coordinates and $g_{ab}$ is a $4\times 4$ symmetric matrix. Physical parameters (bubble velocity $v_{s}$, size $R$, thickness $\sigma$) are dynamic fields that can be modified without recompilation. Grid evaluation lifts the pointwise function to arbitrary grids via jax.vmap, with optional chunked batching via jax.lax.map for memory-limited hardware.
@@ 4 Results: Energy conditions
We evaluate energy conditions on $50^{3}$ spatial grids for each metric, with the time coordinate fixed at $t=0$ (for constant-velocity bubbles whose metric depends on $x-v_{s}t$, the $t=0$ slice is representative up to spatial translation). The grid extends $\pm 5R$ for metrics with $R=1$ (Alcubierre, Van Den Broeck, Natário) and $\pm 3R$ for large-radius metrics (Rodal, Lentz with $R=100$), centered on the bubble; the WarpShell grid extends $\pm 5R_{2}$. For the Alcubierre metric with $\sigma=8$, the shape function $f(r_{s})=\{\tanh[\sigma(r_{s}+R)]-\tanh[\sigma(r_{s}-R)]\}/[2\tanh(\sigma R)]$ has a 10–90% transition width of $0.27$ coordinate units (Table 2), so the $50^{3}$ grid ($h=10/49\approx 0.20$) places $\sim\!1.3$ cells across the wall. For a monotone $\tanh$ wall $g(r)=[1-\tanh(\sigma(r-R))]/2$, the 10–90% levels correspond to $\tanh=\mp 0.8$, giving $\Delta r=2\,\mathrm{atanh}(0.8)/\sigma\approx 2.197/\sigma$. The convergence study (Section 7) confirms that the minimum NEC margin is resolution-stable (within $\sim\!1$% between $50^{3}$ and $100^{3}$) despite this marginal resolution. Table 1 lists the parameters used for each metric.
Table 1: Per-metric parameters and grid configuration. All grids are $50^{3}$ at $t=0$, centered on the bubble.
Metric $R$ $\sigma$ Domain $\zeta_{\max}$
@@
Table 2: Wall resolution analysis for all metrics ($50^{3}$ grid). Transition width computed as the 10–90% threshold distance of each metric’s shape function ($\Delta r=2\,\mathrm{atanh}(0.8)/\sigma$ for $\tanh$-based walls). Autodiff computes exact curvature at each grid point; wall resolution affects spatial sampling density, not curvature accuracy.
Metric Wall width $\Delta x$ Cells Resolved
Alcubierre 0.27 0.20 1.35 No
@@
Rodal 73.24 12.24 5.98 Yes
WarpShell 0.06 0.20 0.29 No
@@
For each grid point, we compute: (i) the Eulerian baseline (ADM normal observer for WEC/SEC/DEC; six axis-aligned null directions for NEC), and (ii) the hybrid margins (algebraic slack at Type I points, $\zeta_{\max}$-capped observer extremum at non-Type I) via multi-start optimization (Section 2.4) with $N_{\text{starts}}=8$. The Hawking–Ellis classifier identifies the vast majority of grid points as Type I (four real eigenvalues with a timelike eigenvector): Rodal achieves 100%, Lentz 99.9%, WarpShell 99.6%, VdB 98.4%, Alcubierre 98%, and Natário $\sim\!97.5$%. The minimum across all tested metrics is Natário at 97.5% (Table 8). At Type I points the algebraic eigenvalue inequalities (equations 1013) apply; at non-Type I points the optimizer provides capped extrema as the sole diagnostic.
@@ 4.1 Individual metric analysis
The original Alcubierre metric [1] with $\tanh$-based shaping function $f(r_{s})$ (equation 1) at $v_{s}=0.5$, $R=1$, $\sigma=8$ is the primary test case. Figure 1 shows the NEC comparison between Eulerian and observer-robust evaluation. The Eulerian analysis identifies NEC violations in an annular region at the bubble wall where the energy density is most negative. The observer-robust analysis detects no additional NEC violation points beyond the Eulerian result (Table 3): the violation sets coincide. However, the violation severity differs: the minimum NEC margin across the grid is 0.626-0.626 (robust) versus 0.625-0.625 (Eulerian), a modest difference for the NEC. The WEC tells a more dramatic story (Figure 2): while the violation sets again coincide (0% missed), the minimum WEC margin at $\zeta_{\max}=5$ ($\gamma\approx 74$) is $\sim\!-3450$ for the optimized observer, compared to 0.038-0.038 for the Eulerian frame. This large ratio arises because at NEC-violating points the WEC energy density $T_{ab}u^{a}u^{b}$ scales as $\sim\!\gamma^{2}$ for highly boosted observers (Section 2.4, Interpretation caveat). Nevertheless, this demonstrates that even when the Eulerian frame correctly identifies where violations occur, boosted observers can experience much larger violation magnitudes.
@@
The Lentz metric [17] was originally proposed as a “positive-energy” warp drive. However, Celmaster and Rubin [21] showed that the Lentz metric violates the WEC even for Eulerian observers once the metric is properly evaluated. The diamond-pattern shift profile uses $\tanh$-based smooth transitions following the WarpFactory parameterization. Like the Alcubierre case, the observer-robust analysis finds no additional NEC violations beyond the Eulerian result (Table 3): the six-direction NEC baseline detects the full NEC-violating region at the tested parameters. The NEC violations detected by both analyses are concentrated near the bubble boundary, consistent with the findings of Celmaster and Rubin. Because the Lentz wall spans only $\sim\!0.02$ grid cells at $50^{3}$ (Table 2), the wall is effectively unresolved and any quantitative Lentz results should be regarded as qualitative indicators rather than converged values.
@@
The Van Den Broeck metric modifies Alcubierre’s construction by introducing a conformal factor $B(r_{s})$ on the spatial metric, expanding the bubble’s internal volume while keeping the external surface area small. Unlike Alcubierre and Lentz, the observer-robust analysis discovers missed NEC, WEC, SEC, and DEC violations (Table 3), with SEC showing the largest conditional miss rate ($f_{\text{miss}|\text{viol}}=17.5\%$; Figure 4). The NEC comparison (Figure 3) confirms that even small miss fractions ($0.1\%$ of grid points) represent genuine additional violations that the Eulerian frame fails to detect.
@@
For each grid point, the WEC optimizer returns the worst-case timelike observer parameters $(\zeta^{*},\theta^{*},\phi^{*})$. Figure 5 visualizes this as a boost direction field overlaid on the WEC margin for the Alcubierre metric. The arrows show the spatial direction and magnitude ($|\sinh\zeta^{*}|$) of the Lorentz boost that minimizes the WEC margin. The worst-case observers are predominantly boosted along the direction of bubble propagation, with larger rapidity values near the bubble wall. This is physically intuitive: an observer boosted into the bubble wall “sees” the negative energy density more prominently.
@@ 4.2 WarpFactory comparison
i.e. the fraction of all grid points where the Eulerian margin is non-negative (condition appears satisfied) but the robust ground truth is negative (condition violated). For Type I points (the dominant case; see Table 8), the ground truth $m_{\text{rob}}$ is the algebraic slack (Section 2.2), which is cap-independent; consequently, $f_{\text{miss}}$ is insensitive to $\zeta_{\max}$ in practice. Both $f_{\text{miss}}$ and “% violated” are domain-restricted diagnostics: their values depend on the sampled region, grid density, and—crucially—the wall thickness of each metric. Metrics with broader walls (e.g. Rodal, $\sigma=0.03$, wall width $\approx 73$ cells; Table 2) naturally occupy more grid volume, inflating volume-based violation fractions relative to thin-wall metrics. These fractions are therefore not directly comparable across metrics as a severity measure; they are primarily meaningful for comparing the Eulerian and robust analyses within the same domain and metric. The conditional miss rate $f_{\text{miss}|\text{viol}}\equiv(\text{missed})/(\text{total violated})$, which normalizes by the number of actually-violated points, is the more appropriate cross-metric diagnostic for our central claim. Table 3 summarizes $f_{\text{miss}}$ for all metrics at $v_{s}=0.5$. The Rodal metric shows substantial missed violations across NEC, WEC, SEC, and DEC. SEC misses are notably widespread: Van Den Broeck has $1.2\%$ SEC missed despite nonzero NEC/WEC/DEC misses, demonstrating that SEC violations are genuinely independent of the other conditions. This independence arises because SEC involves the trace $\rho+\sum_{i}p_{i}$, which can be negative even when each individual $\rho+p_{i}\geq 0$ (NEC satisfied). The violation severity can differ substantially between Eulerian and optimized observers even when the violation sets coincide. Table values are percentages over the full $50^{3}$ volume rounded to $0.1\%$; slice plots may show a few points even when the rounded volume fraction is $0.0\%$. In unrounded values, WarpShell shows NEC total violated at $0.259\%$ and WEC total violated at $0.270\%$. A slight WEC >> NEC difference is expected because the WEC is strictly stronger than the NEC: WEC satisfaction implies NEC satisfaction, but WEC can be violated even when NEC holds. At this sub-percent level, exact counts can also be influenced by finite-precision effects near the violation boundary and by the fact that our non-Type I diagnostics are $\zeta_{\max}$-capped. At the 0.1% reporting precision of Tables 3 and 4, both values round to the same figure; the underlying sub-percent difference is visible only at higher precision. Table 4 shows the velocity dependence for the two metrics with the largest discrepancies.

### 27 Apr 2026 v3

@@ 2.4 Observer optimization
where $m_{\text{flux}}=-g_{ab}\,j^{a}j^{b}$ (positive when the energy flux $j^{a}=-T^{a}{}_{b}\,u^{b}$ is causal), $m_{\text{future}}=-j_{a}n^{a}$ (positive when $j^{a}$ is future-directed, with $n^{a}$ the future-pointing unit normal), and $m_{\text{WEC}}=T_{ab}\,u^{a}u^{b}$ (positive when the energy density is non-negative). A negative value of any component indicates a DEC violation for that observer. The three-term minimum in equation (17) is piecewise smooth but not $C^{1}$ at branch-switching boundaries; gradients exist almost everywhere, and the multi-start BFGS optimizer handles this reliably in practice. A log-sum-exp soft-minimum approximation could be employed if smoother gradients are needed. An alternative design would run three separate optimizations (one per DEC sub-condition) and take the minimum of the minima; this would avoid the unit-mixing objective surface and improve interpretability of the “worst-case observer” at some additional computational cost. Because $m_{\text{WEC}}$ has dimensions of energy density while $m_{\text{flux}}$ has dimensions of (energy density)2, we do not combine them into a single margin for severity purposes. For violation detection, the DEC is violated whenever $m_{\text{DEC}}^{*}<0$ (equation 17); the sign logic is correct regardless of units. For Type I stress-energy with $\rho\geq 0$ and $\rho\geq|p_{i}|$, the algebraic condition (equation 13) guarantees that $j^{a}$ is causal and future-directed for every future-directed timelike $u^{a}$, making the three-term check redundant; the algebraic slack $\min_{i}(\rho-|p_{i}|)$ then serves as the primary DEC result with homogeneous dimensions (energy density). At non-Type I points, the optimizer evaluates all three terms of equation (17) simultaneously (Section 8). For the majority of grid points, which classify as Type I, the algebraic slack is the primary DEC result; at non-Type I points the capped extremum is used instead.
@@
The optimization uses Broyden–Fletcher–Goldfarb–Shanno (BFGS) [33] implemented in Optimistix [24], a pure-JAX optimization library. For timelike observers (WEC/SEC/DEC), the optimizer works over an unconstrained 3-vector $\mathbf{w}\in\mathbb{R}^{3}$, with the rapidity $\zeta=|\mathbf{w}|$ and boost direction $\hat{s}=\mathbf{w}/|\mathbf{w}|$. This parameterization naturally includes the Eulerian observer at $\mathbf{w}=0$ (the removable singularity $\mathbf{w}/|\mathbf{w}|$ at the origin is handled via the smooth approximation $|\mathbf{w}|_{\varepsilon}=\sqrt{\mathbf{w}\cdot\mathbf{w}+\varepsilon^{2}}$ with $\varepsilon=10^{-12}$, which is $C^{\infty}$ everywhere including $\mathbf{w}=0$ and avoids discontinuous gradients that could confuse quasi-Newton Hessian updates near the Eulerian point; see Section 8). This parameterization avoids angular coordinate singularities and ensures a smooth optimization surface. A soft cap $|\mathbf{w}|\mapsto\zeta_{\max}\tanh(|\mathbf{w}|/\zeta_{\max})$ bounds the rapidity at $\zeta_{\max}=5$ (corresponding to $\gamma=\cosh 5\approx 74$); the $\tanh$ map is chosen because it smoothly saturates with unit gradient near the origin, preserving optimizer sensitivity for moderate boosts while enforcing a strict upper bound. For the NEC, the null direction is parameterized via stereographic projection: a 2-vector $\mathbf{w}\in\mathbb{R}^{2}$ maps to a direction on $S^{2}$ via $\hat{s}=(2w_{1},\,2w_{2},\,1-|\mathbf{w}|^{2})/(1+|\mathbf{w}|^{2})$, then $k^{a}=e_{0}^{a}+\hat{s}^{i}\,e_{i}^{a}$. This covers all of $S^{2}$ smoothly without polar coordinate singularities ($\mathbf{w}=0\to$ north pole; $|\mathbf{w}|\to\infty\to$ south pole). Because the south pole requires $|\mathbf{w}|\to\infty$, optimizer convergence near it can be slow; the multi-start pool includes near-pole initializations (axis-aligned directions with large $|\mathbf{w}|$) to mitigate this.
@@ 2.5 Geodesic integration
which is integrated simultaneously with the geodesic equation as a coupled 16-component ordinary differential equation (ODE) system $y=(x^{\mu},v^{\mu},\xi^{\mu},w^{\mu})$, where $v^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}\lambda$ and $w^{\mu}=D\xi^{\mu}/\mathrm{d}\lambda$. The explicit first-order form is
@@ 3 The warpax toolkit
warpax is a Python package built on JAX [7] for hardware-accelerated numerical computation with automatic differentiation. We organize the toolkit into six computational modules:
@@
Each metric is an Equinox module implementing a pointwise signature $x^{\mu}\mapsto g_{ab}(x)$, where $x^{\mu}$ is a 4-vector of coordinates and $g_{ab}$ is a $4\times 4$ symmetric matrix. Physical parameters (bubble velocity $v_{s}$, bubble radius $R_{b}$, wall thickness $\sigma$) are dynamic fields that can be modified without recompilation. Grid evaluation lifts the pointwise function to arbitrary grids via jax.vmap, with optional chunked batching via jax.lax.map for memory-limited hardware.
@@ 4 Results: Energy conditions
We evaluate energy conditions on $50^{3}$ spatial grids for each metric, with the time coordinate fixed at $t=0$ (for constant-velocity bubbles whose metric depends on $x-v_{s}t$, the $t=0$ slice is representative up to spatial translation). The grid extends $\pm 5R_{b}$ for metrics with $R_{b}=1$ (Alcubierre, Van Den Broeck, Natário) and $\pm 3R_{b}$ for large-radius metrics (Rodal, Lentz with $R_{b}=100$), centered on the bubble; the WarpShell grid extends $\pm 5R_{2}$. For the Alcubierre metric with $\sigma=8$, the shape function $f(r_{s})=\{\tanh[\sigma(r_{s}+R_{b})]-\tanh[\sigma(r_{s}-R_{b})]\}/[2\tanh(\sigma R_{b})]$ has a 10–90% transition width of $0.27$ coordinate units (Table 3), so the $50^{3}$ grid ($h=10/49\approx 0.20$) places $\sim\!1.3$ cells across the wall. For a monotone $\tanh$ wall $g(r)=[1-\tanh(\sigma(r-R_{b}))]/2$, the 10–90% levels correspond to $\tanh=\mp 0.8$, giving $\Delta r=2\,\mathrm{atanh}(0.8)/\sigma\approx 2.197/\sigma$. The convergence study (Section 7) confirms that the minimum NEC margin is resolution-stable (within $\sim\!1$% between $50^{3}$ and $100^{3}$) despite this marginal resolution. Table 1 lists the parameters used for each metric.
Table 1: Per-metric parameters and grid configuration. All grids are $50^{3}$ at $t=0$, centered on the bubble. Matched-parameter cross-metric comparison ($R_{b}=1$, $\sigma=8$) was attempted for Rodal; NEC $f_{\text{miss}}$ is resolution-unstable at these parameters (Table 2), so Rodal is reported at its native $R_{b}=100$, $\sigma=0.03$.
Metric $R_{b}$ $\sigma$ Domain $\zeta_{\max}$
@@
Table 2: Matched-parameter diagnostic: Rodal at $R_{b}=1$, $\sigma=8$ on domain $(\pm 3)^{3}$. The NEC missed-violation fraction is resolution-unstable (max deviation ${\sim}9.9\%$ across $N\in\{30,50,70\}$), exceeding the $\pm 5\%$ stability criterion. Reported as a domain-restricted diagnostic; the main comparison (Table 1) uses Rodal at native $R_{b}=100$, $\sigma=0.03$.
NN $n_{\text{total}}$ NEC $f_{\text{miss}}$ (%) Stability
30 27,000 0.88 Unstable
50 125,000 1.01
70 343,000 1.06
Table 3: Wall resolution analysis for all metrics ($50^{3}$ grid). Transition width computed as the 10–90% threshold distance of each metric’s shape function ($\Delta r=2\,\mathrm{atanh}(0.8)/\sigma$ for $\tanh$-based walls). The Convergence column records the resolution-support tier: Richardson (numerical extrapolation), Stability-only (minimum margin stable under refinement), Weakest (least refined), – (no convergence study). Autodiff computes exact curvature at each grid point; wall resolution affects spatial sampling density, not curvature accuracy.
Metric Wall width $\Delta x$ Cells Resolved Convergence
Alcubierre 0.27 0.20 1.35 No Richardson
Natário 0.27 0.20 1.35 No
@@
WarpShell 0.06 0.20 0.29 No Weakest
Schwarzschild 0.82 Richardson
@@
For each grid point, we compute: (i) the Eulerian baseline (ADM normal observer for WEC/SEC/DEC; six axis-aligned null directions for NEC), and (ii) the hybrid margins (algebraic slack at Type I points, $\zeta_{\max}$-capped observer extremum at non-Type I) via multi-start optimization (Section 2.4) with $N_{\text{starts}}=8$. Our Hawking–Ellis classifier identifies the vast majority of grid points as Type I (four real eigenvalues with a timelike eigenvector): Rodal achieves 100%, Lentz 99.9%, WarpShell 99.6%, VdB 98.4%, Alcubierre 98%, and Natário $\sim\!97.5$%. The minimum across all tested metrics is Natário at 97.5% (Table 9). At Type I points the algebraic eigenvalue inequalities (equations 1013) apply; at non-Type I points the optimizer provides capped extrema as the sole diagnostic.
@@ 4.1 Individual metric analysis
The original Alcubierre metric [1] with $\tanh$-based shaping function $f(r_{s})$ (equation 1) at $v_{s}=0.5$, $R_{b}=1$, $\sigma=8$ serves as our primary test case. Figure 1 shows the NEC comparison between Eulerian and observer-robust evaluation. The Eulerian analysis identifies NEC violations in an annular region at the bubble wall where the Eulerian energy density $T_{ab}\,n^{a}n^{b}$ is most negative. The observer-robust analysis finds no additional NEC violation points beyond the Eulerian result (Table 4): the violation sets coincide. However, the violation severity differs: the minimum NEC margin across the grid is 0.626-0.626 (robust) versus 0.625-0.625 (Eulerian), a modest difference for the NEC. The WEC tells a more striking story (Figure 2): while the violation sets again coincide (0% missed), the minimum WEC margin at $\zeta_{\max}=5$ ($\gamma\approx 74$) is $\sim\!-3450$ for the optimized observer, compared to 0.038-0.038 for the Eulerian frame. This large ratio arises because at NEC-violating points the WEC energy density $T_{ab}u^{a}u^{b}$ scales as $\sim\!\gamma^{2}$ for highly boosted observers (Section 2.4, Interpretation caveat). Nevertheless, this demonstrates that even when the Eulerian frame correctly identifies where violations occur, boosted observers can experience much larger violation magnitudes.
@@
The Lentz metric [25] was originally proposed as a “positive-energy” warp drive. However, Celmaster and Rubin [8] showed that it violates the WEC even for Eulerian observers once the metric is properly evaluated. The diamond-pattern shift profile uses $\tanh$-based smooth transitions following the WarpFactory parameterization. Like the Alcubierre case, the observer-robust analysis finds no additional NEC violations beyond the Eulerian result (Table 4): the six-direction NEC baseline detects the full NEC-violating region at the tested parameters. The NEC violations detected by both analyses are concentrated near the bubble boundary, consistent with the findings of Celmaster and Rubin. Because the Lentz wall spans only $\sim\!0.02$ grid cells at $50^{3}$ (Table 3), the wall is effectively unresolved and any quantitative Lentz results should be regarded as qualitative indicators rather than converged values.
@@
The Van Den Broeck metric modifies Alcubierre’s construction by introducing a conformal factor $B(r_{s})$ on the spatial metric, expanding the bubble’s internal volume while keeping the external surface area small. Unlike Alcubierre and Lentz, our observer-robust analysis discovers missed NEC, WEC, SEC, and DEC violations (Table 4), with SEC showing the largest conditional miss rate ($f_{\text{miss}|\text{viol}}=17.5\%$; Figure 4). The NEC comparison (Figure 3) confirms that even small miss fractions ($0.1\%$ of grid points) represent genuine additional violations that the Eulerian frame fails to detect.
@@
For each grid point, the WEC optimizer returns the worst-case timelike observer parameters $(\zeta^{*},\theta^{*},\phi^{*})$. Figure 5 visualizes this as a boost direction field overlaid on the WEC margin for the Alcubierre metric. The arrows show the spatial direction and magnitude ($|\sinh\zeta^{*}|$) of the Lorentz boost that minimizes the WEC margin. The worst-case observers are predominantly boosted along the direction of bubble propagation, with larger rapidity values near the bubble wall. This is physically intuitive: an observer boosted into the bubble wall “sees” a more negative observer-contracted energy density $T_{ab}\,u^{a}u^{b}$ than an unboosted Eulerian observer at the same spacetime point.
@@ 4.2 WarpFactory comparison
i.e. the fraction of all grid points where the Eulerian margin is non-negative (condition appears satisfied) but the robust ground truth is negative (condition violated). For Type I points (the dominant case; see Table 9), the ground truth $m_{\text{rob}}$ is the algebraic slack (Section 2.2), which is cap-independent; consequently, $f_{\text{miss}}$ is insensitive to $\zeta_{\max}$ in practice. Both $f_{\text{miss}}$ and “% violated” are domain-restricted diagnostics: their values depend on the sampled region, grid density, and—crucially—the wall thickness of each metric. Metrics with broader walls (e.g. Rodal, $\sigma=0.03$, wall width $\approx 73$ cells; Table 3) naturally occupy more grid volume, inflating volume-based violation fractions relative to thin-wall metrics. These fractions are therefore not directly comparable across metrics as a severity measure; they are primarily meaningful for comparing the Eulerian and robust analyses within the same domain and metric. The conditional miss rate $f_{\text{miss}|\text{viol}}\equiv(\text{missed})/(\text{total violated})$, which normalizes by the number of actually-violated points, is the more appropriate cross-metric diagnostic for our central claim. Table 4 summarizes $f_{\text{miss}}$ for all metrics at $v_{s}=0.5$. The Rodal metric shows substantial missed violations across NEC, WEC, SEC, and DEC. SEC misses are notably widespread: Van Den Broeck has $1.2\%$ SEC missed despite nonzero NEC/WEC/DEC misses, demonstrating that SEC violations are genuinely independent of the other conditions. This independence arises because SEC involves the trace $\rho+\sum_{i}p_{i}$, which can be negative even when each individual $\rho+p_{i}\geq 0$ (NEC satisfied). The violation severity can differ substantially between Eulerian and optimized observers even when the violation sets coincide. Table values are percentages over the full $50^{3}$ volume rounded to $0.1\%$; slice plots may show a few points even when the rounded volume fraction is $0.0\%$. In unrounded values, WarpShell shows NEC total violated at $0.259\%$ and WEC total violated at $0.270\%$. A slight WEC >> NEC difference is expected because the WEC is strictly stronger than the NEC: WEC satisfaction implies NEC satisfaction, but WEC can be violated even when NEC holds. At this sub-percent level, exact counts can also be influenced by finite-precision effects near the violation boundary and by the fact that our non-Type I diagnostics are $\zeta_{\max}$-capped. At the 0.1% reporting precision of Tables 4 and 5, both values round to the same figure; the underlying sub-percent difference is visible only at higher precision. Table 5 shows the velocity dependence for the two metrics with the largest discrepancies.
@@
 Lentz wall is severely under-resolved at $50^{3}$ ($\sim\!0.02$ cells; Table 3); fractions are unresolved lower-bound diagnostics, not directly comparable to better-resolved metrics.
 WarpShell uses a $C^{2}$ quintic Hermite regularization of the thin-shell; margins reflect the regularized implementation, not an idealized thin-shell spacetime.
@@
 WarpShell uses a $C^{2}$ quintic Hermite regularization of the thin-shell; margins reflect the regularized implementation, not an idealized thin-shell spacetime.